Required information [The following information applies to the questions displayed below.] Forten Companys current year incoFORTEN COMPANY Income Statement For Current Year Ended December 31 Sales Cost of goods sold Gross profit Operating expenses DRequired: 1. Prepare a complete statement of cash flows using the indirect method for the current year. (Amounts to be deductFORTEN COMPANY Statement of Cash Flows For Current Year Ended December 31 Cash flows from operating activities Adjustments toRequired: Prepare a complete statement of cash flows using a spreadsheet using the indirect method. (Enter all amounts as posFORTEN COMPANY Spreadsheet for Statement of Cash Flows For Current Year Ended December 31 Anal ysis of Changes December 31 De
Answer #1
Cash Flow Statement
Indirect Method
Cash Flow from Operating Activities
Net Income $        107,375
Adjustments
Depreciation $        39,750
Loss on Sale of Equipment $        24,125
Changes in Current Assets/ Current Liabilities
Increase in Accounts Receivable $      -24,835 =69625-94460
Increase in inventory $      -33,356 =270800-304156
Decrease in Prepaid Rent $              875 =2275-1400
Decrease in Accounts Payable $      -71,034 =72141-143175
Total Adjustments $        -64,475
Cash from Operating Activities $         42,900
Cash flow from Investing Activities
Cash paid for Purchase of Equipment $        -68,000
Sale of Equipment $          30,625
Net cash used In investing activities $       -37,375
Cash flow from Financing Activities
Issue of Common Stock $          88,000 =4400*20
Borrowed Loan by Short term note $            5,900
Repayment of Long term notes $        -59,625
Dividend Paid $        -53,900
Net cash used in financing activities $       -19,625
Increase in Cash $       -14,100
Opening Balance of Cash $         92,500
Closing Balance of Cash $         78,400
Dec 31, Prior Year Debit Credit Dec 31, Current Year
Balance Sheet-Debit
Cash $       92,500 $         14,100 $           78,400
Accounts Receivable $       69,625 $          24,835 $           94,460
Inventory $     270,800 $          33,356 $         304,156
Prepaid Expenses $          2,275 $               875 $             1,400
Equipment $     127,000 $        115,375 $       103,875 $         138,500
$     562,200 $         616,916
Balance Sheet-Credit
Accumulated Depreciation $       55,500 $          49,125 $         39,750 $           46,125
Accounts Payable $     143,175 $          71,034 $           72,141
Short term notes payable $          9,800 $           5,900 $           15,700
Long term notes payable $       67,750 $          59,625 $         47,375 $           55,500
Common Stock $     169,250 $         22,000 $         191,250
Paid in Capital in excess of par value $                 -   $         66,000 $           66,000
Retained Earnings $     116,725 $          53,900 $       107,375 $         170,200
$     562,200 $         616,916